OpenStudy (anonymous):

I need help with understanding the substitution part! thanks Definite Integral w/ substitution 2t^2(1-4t^3) dt with the limits x=0 and x=-2

myininaya (myininaya):

\[\int\limits_{0}^{2}2t^2(1-4t^3) dt \] So we have something like this: \[\int\limits_{a}^{b}c f'(t) f(t) dt \] u=f(t) du/dt=f'(t) du=f'(t) dt If t=a, u=f(a) If t=b, u=f(b) \[\int\limits_{f(a)}^{f(b)}c u du =c \frac{u^2}{2}|_{f(a)}^{f(b)}=\frac{c}{2}((f(b))^2-f(a))^2)\]

OpenStudy (anonymous):

thank you... the part I'm having a hard time with is the substitution... I know that i would state that u= 1-4t^3 and du = 12t^2 ??? so du/dt equals???

myininaya (myininaya):

du=-12t^2 dt

myininaya (myininaya):

\[-1 \int\limits_{-2}^{0} -1 \cdot 2t^2(1-4t^3) dt \] By the by -1(-1)=1 So we have